# Tackling Long-Horizon Tasks with Model-based Offline Reinforcement Learning

### v1

@@ 5.4 Results on MuJoCo Gym locomotion tasks
Dataset BC TD3+BC CQL EDAC IQL MOPO COMBO RAMBO MOBILE CBOP LEQ (ours)
hopper-r $3.7$ $8.5$ $5.3$ $25.3$ $7.6$ $31.7$ $17.9$ $25.4$ $\mathbf{31.9}$ $\mathbf{32.8}$ $\mathbf{32.4}$$\pm 0.3$
hopper-m $54.1$ $59.3$ $61.9$ $\mathbf{101.6}$ $66.3$ $62.8$ $97.2$ $87.0$ $\mathbf{106.6}$ $\mathbf{102.6}$ $\mathbf{103.4}$$\pm 0.3$
hopper-mr $16.6$ $60.9$ $86.3$ $\mathbf{101.0}$ $94.7$ $\mathbf{99.4}$ $\mathbf{103.5}$ $89.5$ $99.5$ $\mathbf{104.3}$ $\mathbf{103.9}$$\pm 1.3$
hopper-me $53.9$ $98.0$ $96.9$ $\mathbf{110.7}$ $91.5$ $81.6$ $\mathbf{111.1}$ $88.2$ $\mathbf{112.6}$ $\mathbf{111.6}$ $\mathbf{109.4}$$\pm 1.8$
walker2d-r $1.3$ $1.6$ $5.4$ $16.6$ $5.2$ $7.4$ $7.0$ $0.0$ $17.9$ $17.8$ $\mathbf{21.5}$$\pm 0.1$
walker2d-m $70.9$ $83.7$ $79.5$ $\mathbf{92.5}$ $78.3$ $81.3$ $84.1$ $81.9$ $84.9$ $\mathbf{87.7}$ $74.9$$\pm 26.9$
walker2d-mr $20.3$ $81.8$ $76.8$ $87.1$ $73.9$ $85.6$ $56.0$ $89.2$ $89.9$ $92.7$ $\mathbf{98.7}$$\pm 6.0$
walker2d-me $90.1$ $110.1$ $109.1$ $\mathbf{114.7}$ $109.6$ $\mathbf{112.9}$ $103.3$ $56.7$ $\mathbf{115.2}$ $\mathbf{117.2}$ $108.2$$\pm 1.3$
halfcheetah-r $2.2$ $11.0$ $31.3$ $28.4$ $11.8$ $\mathbf{38.5}$ $\mathbf{38.8}$ $\mathbf{39.5}$ $\mathbf{39.3}$ $32.8$ $30.8$$\pm 3.3$
halfcheetah-m $43.2$ $48.3$ $46.9$ $65.9$ $47.4$ $73.0$ $54.2$ $\mathbf{77.9}$ $\mathbf{74.6}$ $\mathbf{74.3}$ $71.7$$\pm 4.4$
halfcheetah-mr $37.6$ $44.6$ $45.3$ $61.3$ $44.2$ $\mathbf{72.1}$ $55.1$ $\mathbf{68.7}$ $\mathbf{71.7}$ $66.4$ $65.5$$\pm 1.1$
halfcheetah-me $44.0$ $90.7$ $95.0$ $\mathbf{106.3}$ $86.7$ $90.8$ $90.0$ $95.4$ $\mathbf{108.2}$ $\mathbf{105.4}$ $\mathbf{102.8}$$\pm 0.4$
Total $437.9$ $698.5$ $739.7$ $911.4$ $717.2$ $844.0$ $802.0$ $812.4$ $\mathbf{959.5}$ $\mathbf{953.4}$ $\mathbf{923.2}$
@@ 5.5 Ablation studies
To understand why LEQ (LEQ-$\lambda$) works well in long-horizon tasks, we conduct ablation studies and answer to the following four questions: (1) Does using $\lambda$-return help? (2) Is LEQ better than prior uncertainty-based penalization methods? (3) Which factor enables LEQ to work in AntMaze? and (4) How do imagination length $H$ and data expansion length $R$ affect the performance?
@@
To verify the effect of $\lambda$-return, we compare our method (LEQ-$\lambda$) with the versions with $1$-step return (LEQ-$1$) and $H$-step return (LEQ-$H$). Table 4 shows that using $\lambda$-return drastically improves the performance on AntMaze compared to using $1$-step return or $H$-step return. This result is coherent with the observations in prior online RL methods [28, 11].
(2) Lower expectile Q-learning.
We compare our lower expectile Q-learning with another conservative value estimator, MOBIP used in MOBILE [30], which penalizes Q-values with the standard deviation of Q-ensemble networks. The only difference between LEQ and MOBIP is their target Q-value computation for both critic and policy updates. Table 4 shows that using MOBIP not only deteriorates the success rates (in MOBIP-$1$) but also does not benefit from $\lambda$-return (in MOBIP-$\lambda$).
(3) What makes offline model-based RL work in AntMaze?
Prior to LEQ, none of offline model-based RL methods work in AntMaze, whereas our method even outperforms model-free methods. Thus, we investigate which changes in LEQ enable offline model-based RL work in AntMaze.
Table 4: Ablation study results on the AntMaze tasks. (1) We compare different Q-targets, LEQ-$\lambda$, LEQ-$1$, and LEQ-$H$. (2) We compare our lower expectile Q-learning strategy with another conservative Q-value estimation, MOBIP-$1$ and MOBIP-$\lambda$. (3) MOBILE is our re-implementation of MOBILE using MOBILE’s default hyperparameters, i.e., $\beta=0.95$, $\gamma=0.99$, and $R=5$. We note that lowering $\beta$ to $0.25$ is crucial for MOBILE to achieve meaningful scores in AntMaze.
Dataset umaze medium large ultra Total
umaze diverse play diverse play diverse play diverse
LEQ-$\lambda$ (ours) $\mathbf{94.4}$$\pm 6.3$ $\mathbf{71.0}$$\pm 12.3$ $58.8$$\pm 33.0$ $46.2$$\pm 23.2$ $\mathbf{58.6}$$\pm 9.1$ $\mathbf{60.2}$$\pm 18.3$ $25.8$$\pm 18.2$ $\mathbf{55.8}$$\pm 18.3$ $\mathbf{470.4}$
LEQ-$H$ $\mathbf{93.0}$$\pm 3.4$ $60.7$$\pm 10.4$ $46.3$$\pm 32.4$ $0.0$ ±$\pm 0.0$ $\mathbf{57.0}$$\pm 25.6$ $33.3$$\pm 43.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $290.3$
LEQ-$1$ $89.6$$\pm 4.8$ $37.0$$\pm 32.8$ $55.8$$\pm 28.7$ $29.8$$\pm 24.5$ $34.2$$\pm 13.4$ $49.3$$\pm 9.0$ $\mathbf{42.2}$$\pm 13.2$ $35.6$$\pm 13.0$ $373.5$
MOBIP-$\lambda$ $84.3$$\pm 3.5$ $40.3$$\pm 20.4$ $51.3$$\pm 9.0$ $39.7$$\pm 12.5$ $28.3$$\pm 21.5$ $33.7$$\pm 10.0$ $38.0$$\pm 27.1$ $23.3$$\pm 4.9$ $338.9$
MOBIP-1 $59.5$$\pm 3.5$ $46.5$$\pm 1.5$ $57.0$$\pm 11.0$ $\mathbf{54.0}$$\pm 9.0$ $23.5$$\pm 19.5$ $38.5$$\pm 1.5$ $39.5$$\pm 11.5$ $20.5$$\pm 20.5$ $339.0$
MOBILE $1.0$$\pm 2.0$ $0.0$ ±$\pm 0.0$ $6.4$$\pm 5.5$ $5.0$$\pm 5.0$ $0.8$$\pm 1.6$ $0.8$$\pm 1.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $14.0$
MOBILE ($\mathbf{\beta}$ = 0.25) $77.0$$\pm 6.4$ $20.4$$\pm 15.7$ $\mathbf{64.6}$$\pm 11.1$ $31.6$$\pm 16.9$ $2.6$$\pm 2.8$ $7.2$$\pm 8.9$ $4.6$$\pm 3.0$ $5.0$$\pm 4.6$ $213.0$
MOBILE ($\mathbf{\gamma}$ = 0.997) $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $7.2$$\pm 4.1$ $1.6$$\pm 2.1$ $9.6$$\pm 7.1$ $5.4$$\pm 4.9$ $0.0$ ±$\pm 0.0$ $1.8$$\pm 2.7$ $25.6$
MOBILE ($R$ = 10) $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $5.0$$\pm 5.1$ $0.6$$\pm 1.2$ $7.4$$\pm 14.8$ $1.6$$\pm 3.2$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $14.6$
@@
We first re-implement MOBILE with some technical tricks used in LEQ: LayerNorm [3], SymLog [11], single Q-network, and no target Q-value clipping; but, MOBILE achieves a barely non-zero score, $14.0$. We found that the key to make MOBILE work is reducing $\beta$, the ratio for the loss calculated from imaginary rollouts and from dataset transitions. When we lower $\beta$ from $0.95$ to $0.25$ (used in LEQ), MOBILE shows meaningful performances in umaze and medium mazes, and achieves $213.0$ in total. We suggest that utilizing the true transition from the dataset is important in long-horizon tasks, which was undervalued in prior works.
(4) Imagination length $H$ and dataset expansion length $R$.
@@
As shown in Table 5, the performance increases when it goes to $H=10$ from $H=5$, but it drops when $H=15$. This result shows the trade-off of using the world model: the further the agent imagines, more the agent becomes robust to the error of the critic, but more it becomes prone to the error from the model prediction.
@@
We also evaluate LEQ without the dataset expansion ($H=10,R=1$). In AntMaze, the results with and without the dataset expansion are similar, as shown in Table 5. On the other hand, the dataset expansion makes the policy more stable and better in the D4RL MuJoCo tasks (Table 13).
@@
Table 5: LEQ with different imagination length $H$ and data expansion length $R$. A longer $H$ can mitigate critic biases, while increasing model errors, which leads to poor performance. Each number is averaged over $5$ random seeds.
Dataset $\mathbf{H=10,R=5}$ (ours) $H=5,R=5$ $H=15,R=5$ $H=10,R=1$
@@ 6 Conclusion
In this paper, we propose a novel offline model-based reinforcement learning method, LEQ, which uses expectile regression to get a conservative evaluation of a policy from model-generated trajectories. Expectile regression eases the pain of constructing the whole distribution of Q-targets and allows for estimating the conservative value via sampling. Combined with $\lambda$-returns in both critic and policy updates for the imaginary rollouts, the policy can receive learning signals that are more robust to both model errors and critic errors. We empirically show that LEQ improves the performance in various tasks – especially, achieving the state-of-the-art performance in the long-horizon AntMaze tasks.
6.1 Limitations
Following prior work on model-based offline RL [30, 14], we assume access to the ground-truth termination function of a task, different from online model-based RL approaches, which learn a termination function from interactions. However, since this termination function is conditioned on a state, a model requires to plan on a state space (or an observation space), which could be challenging in a high-dimensional state space (e.g. pixel observations). Extending the proposed approach to complex environments with high-dimensional observations would be an immediate next step.
6.2 Broader Impacts
Our method aims to increase the ability of autonomous agents, such as robots and self-driving cars, to learn from static, offline data without interacting with the world. This enables autonomous agents to utilize data with diverse qualities (not necessarily from experts). We believe that this paper does not have any immediate negative societal impact.
Acknowledgments and Disclosure of Funding
We would like to thank Junik Bae for helpful discussion. This work was supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant (RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)) and the National Research Foundation of Korea (NRF) grant (RS-2024-00333634) funded by the Korean Government (MSIT). Kwanyoung Park was supported by Electronics and Telecommunications Research Institute (ETRI).
@@ Appendix A Training Details
All experiments are done on a single RTX 4090 GPU and $4$ AMD EPYC 9354 CPU cores. We use $5$ different random seeds for each experiment and report the mean and standard deviation. Each offline RL experiment takes $2$ hours for ours, $12$ hours for MOBILE, and $24$ hours for CBOP.
@@
For the locomotion tasks, we use the dataset provided by D4RL [6] and NeoRL [26]. Following IQL [17], we normalize rewards using the maximum and minimum return of all trajectories. We use the true termination functions of the environments, implemented in MOBILE [30].
For the AntMaze tasks, we use the dataset provided by D4RL [6]. Following IQL [17], we subtract $1$ from the rewards in the datasets so that the agent receives $-1$ for each step and 00 on termination. We use the true termination functions of the environments. The termination functions of the AntMaze tasks are not deterministic because a goal of a maze is randomized every time the environment is reset. Nevertheless, we follow the implementation of CBOP [14], where the termination region is set to a circle around the mean of the goal distribution with the radius $0.5$.
Method implementation details.
For all compared methods, we use the results from their corresponding papers when available. For IQL [17], we run the official implementation with $5$ seeds to reproduce the results for the random datasets in D4RL and NeoRL. For the AntMaze tasks, we run the official implementation of MOBILE and CBOP with $5$ random seeds. Please note that the original MOBILE implementation does not use the true termination function, so we replace it with our termination function. For MOPO, COMBO, and RAMBO, we use the results reported in RAMBO [27].
@@
For training world models, we use the architecture and training script from OfflineRL-Kit [29], matching the implementation of MOBILE [30]. Each world model is implemented as a $4$-layer MLPs with the hidden layer size of $200$. We construct an ensemble of world models by selecting $5$ out of $7$ models with the best validation scores. We pretrain the ensemble of world models for each of $5$ random seeds (i.e. training in total $35$ world models and using $25$ models), which takes approximately $5$ hours in average.
@@
We use $3$-layer MLPs with size of $256$ both for the policy network and the critic network. We use layer normalization [3] to prevent catastrophic over/underestimation [4], and squash the state inputs using symlog to keep training stable from outliers in long-horizon model rollouts [11].
Pretraining.
For some environments, we found that a randomly initialized policy can lead to abnormal rewards or transition prediction from the world models in the early stage, leading to unstable training. Following CBOP [14], we pretrain a policy $\pi_{\theta}$ and a critic $Q_{\phi}$ using behavioral cloning and FQE [20], respectively. We use a slightly different implementation of FQE from the original implementation, where the $\arg\min$ operation is approximated with mini-batch gradient descent, similar to standard Q-learning as shown in Algorithm 2.
Algorithm 2 FQE: Fitted Q Evaluation [20]
@@
4:     $y_{i}=\texttt{sg}(Q_{\phi}(\mathbf{s}^{\prime}_{i},\pi_{\theta}(\mathbf{s}^{\prime}_{i})))$triangleright $\texttt{sg}(\cdot)$ is stop-gradient operator
@@
Table 6: Comparisons with the prior model-based approaches and the baseline methods introduced for our ablation studies. We color the hyperparameters in blue if they are the same with LEQ. Otherwise, we color them in red.
Components CBOP MOBILE MOBILE MOBIP LEQ (ours)
Training scheme MVE [5] MBPO [13] MBPO [13] Dyna [32] Dyna [32]
Conservatism Lower-confidence bound Lower-confidence bound Lower-confidence bound Lower-confidence bound Lower expectile
Policy Stochastic Stochastic Stochastic Deterministic Deterministic
Policy objective $Q(\mathbf{s},\mathbf{a})$ $Q(\mathbf{s},\mathbf{a})$ $Q(\mathbf{s},\mathbf{a})$ $\lambda$-returns $\lambda$-returns
Policy pretraining BC BC BC
# of critics 20-50 2 1 1 1
Critic objective Multi-step (adaptive weighting) One-step One-step $\lambda$-returns $\lambda$-returns
Critic pretraining FQE [20] FQE [20] FQE [20]
Horizon length ($H$) 10 1 1 10 10
Rollout length ($R$) 1 or 5 10 5 5
Discount rate ($\gamma$) 0.99 0.99 0.997 0.997 0.997
$\beta$ in Equation 7 1.0 0.95 0.25 0.25 0.25
Impl. tricks Clip Q-values with 00 LayerNorm + Symlog LayerNorm + Symlog LayerNorm + Symlog
Running time 24h 12h 40m 4h 2h

### v2

@@ 5.4 Results on MuJoCo Gym locomotion tasks
Method BC CQL DrQ+BC ACRO OfflineDV2 MOPO ROSMO LEQ
walker_walk-random $2.0$ $14.4$ $5.5$ $0.0$ $\mathbf{28.7}$ $3.2$$\pm 0.4$ $2.9$$\pm 0.4$ $22.4$$\pm 1.1$
walker_walk-medium $40.9$ $14.8$ $\mathbf{46.8}$ $\mathbf{48.7}$ $34.1$ $37.1$$\pm 3.7$ $\mathbf{49.8}$$\pm 2.3$ $43.1$$\pm 3.2$
walker_walk-medium_replay $16.5$ $11.4$ $28.7$ $27.8$ $\mathbf{56.5}$ $11.4$$\pm 9.2$ $28.1$$\pm 0.9$ $43.0$$\pm 7.3$
walker_walk-medium_expert $47.7$ $56.4$ $\mathbf{86.4}$ $\mathbf{91.4}$ $43.9$ $46.6$$\pm 3.5$ $82.9$$\pm 0.7$ $\mathbf{87.2}$$\pm 2.4$
cheetah_run-random $0.0$ $5.9$ $5.8$ $0.0$ $\mathbf{31.7}$ $3.2$$\pm 3.9$ $2.5$$\pm 0.5$ $14.8$$\pm 1.0$
cheetah_run-medium $\mathbf{51.6}$ $40.9$ $\mathbf{53.0}$ $\mathbf{52.8}$ $17.2$ $34.7$$\pm 6.6$ $\mathbf{50.8}$$\pm 2.5$ $37.9$$\pm 8.0$
cheetah_run-medium_replay $25.0$ $10.7$ $44.8$ $41.7$ $\mathbf{61.5}$ $30.3$$\pm 5.1$ $48.6$$\pm 1.5$ $34.3$$\pm 1.1$
cheetah_run-medium_expert $\mathbf{57.5}$ $20.9$ $50.6$ $46.6$ $10.4$ $36.6$$\pm 5.6$ $45.4$$\pm 2.2$ $25.1$$\pm 6.6$
Total $241.2$ $175.4$ $\mathbf{321.6}$ $\mathbf{309.0}$ $284.0$ $203.3$ $\mathbf{310.9}$ $\mathbf{307.6}$
Similar to antmaze-medium, LEQ experiences high variance in MuJoCo tasks. During training, LEQ often achieves high performance, but then, suddenly falls back to 00, as shown in Figure 6. This is mainly because the learned models sometimes fail to capture failures (e.g. hopper and walker falling off) and predict an optimistic future (e.g. hopper and walker walking forward).
5.5 Results with visual inputs
As shown in Table 4, LEQ combined with DreamerV3 (Hafner et al., 2023) performs on par with the state-of-the-art methods on V-D4RL datasets, demonstrating its scalability to visual observations. Notably, LEQ achieves the highest score on the walker_walk-medium_expert dataset among model-based methods, where OfflineDV2 and MOPO struggles. LEQ also outperforms model-free approaches on random datasets and walker_walk-medium_replay dataset, highlighting the strength of model-based methods in datasets with diverse state-action distributions.
@@ 5.6 Ablation studies
To deeply understand how LEQ (LEQ-$\lambda$) work, we conduct ablation studies in AntMaze environments and answer to the following five questions: (1) Is lower expectile Q-learning better than prior uncertainty-based penalization methods? (2) Does lower expectile policy learning better than existing policy learning methods? (3) Does $\mathbf{\lambda}$-return help? (4) Which factor enables LEQ to work in AntMaze? and (5) How do imagination length $H$ and data expansion affect the performance?
(1) Lower expectile Q-learning.
We compare our lower expectile Q-learning with the conservative value estimator in MOBILE (Sun et al., 2023), which penalizes Q-values based on the standard deviation of Q-ensemble networks. In Table 5, replacing lower expectile Q-learning with MOBILE decreases the success rate, both with $\lambda$-returns ($461.8$ vs $338.9$) and without them ($373.5$ vs $339.0$).
(2) Lower expectile policy learning.
We also compare our lower expectile policy learning with AWR (Peng et al., 2019) and directly maximizing $Q(s,a)$. As shown in Table 5, LEQ shows better performance compared to AWR ($461.8$ vs $114.2$) and maximizing the $Q$-values ($461.8$ vs $410.3$).
@@
The first three rows of Table 5 show the effect of $\lambda$-returns in LEQ. Substituting $\lambda$-returns with $H$-step returns ($461.8$ vs $290.3$) or $1$-step returns ($461.8$ vs $373.5$) significantly decreases the performance. Moreover, while LEQ shows better performance than MOBILE without $\lambda$-returns ($373.5$ vs $339.0$), the performance of LEQ gets significantly better with $\lambda$-returns, compared to MOBILE with $\lambda$-returns ($461.8$ vs $338.9$).
Table 5: Impact of lower expectile Q-learning and $\lambda$-returns on AntMaze. We ablate the effects of lower expectile and $\lambda$-returns on the critic and policy updates in LEQ. The design choices from LEQ are colored in blue and other options are colored in red. The results are averaged over $5$ seeds.
Design choices umaze medium large ultra Total
conservatism critic update policy update umaze diverse play diverse play diverse play diverse
Lower expectile $\lambda$-returns $\lambda$-returns $\mathbf{94.4}$$\pm 6.3$ $\mathbf{71.0}$$\pm 12.3$ $50.2$$\pm 39.9$ $46.2$$\pm 23.2$ $\mathbf{58.6}$$\pm 9.1$ $\mathbf{60.2}$$\pm 18.3$ $25.8$$\pm 18.2$ $\mathbf{55.8}$$\pm 18.3$ $\mathbf{461.8}$
Lower expectile $H$-step $H$-step $\mathbf{93.0}$$\pm 3.4$ $60.7$$\pm 10.4$ $46.3$$\pm 32.4$ $0.0$ ±$\pm 0.0$ $\mathbf{57.0}$$\pm 25.6$ $33.3$$\pm 43.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $290.3$
Lower expectile $1$-step $Q(\mathbf{s},\mathbf{a})$ $89.6$$\pm 4.8$ $37.0$$\pm 32.8$ $55.8$$\pm 28.7$ $29.8$$\pm 24.5$ $34.2$$\pm 13.4$ $49.3$$\pm 9.0$ $\mathbf{42.2}$$\pm 13.2$ $35.6$$\pm 13.0$ $373.5$
Lower expectile $\lambda$-returns $Q(\mathbf{s},\mathbf{a})$ $81.0$$\pm 10.5$ $46.2$$\pm 16.8$ $\mathbf{61.8}$$\pm 12.4$ $40.6$$\pm 11.4$ $39.2$$\pm 12.5$ $40.5$$\pm 11.7$ $\mathbf{42.8}$$\pm 21.8$ $47.5$$\pm 5.9$ $410.3$
Lower expectile $\lambda$-returns AWR $69.2$$\pm 7.5$ $44.4$$\pm 18.4$ $0.6$$\pm 0.6$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $0.0$ ±$\pm 0.0$ $114.2$
MOBILE $\lambda$-returns $\lambda$-returns $84.3$$\pm 3.5$ $40.3$$\pm 20.4$ $51.3$$\pm 9.0$ $39.7$$\pm 12.5$ $28.3$$\pm 21.5$ $33.7$$\pm 10.0$ $38.0$$\pm 27.1$ $23.3$$\pm 4.9$ $338.9$
MOBILE $1$-step $Q(\mathbf{s},\mathbf{a})$ $59.5$$\pm 3.5$ $46.5$$\pm 1.5$ $57.0$$\pm 11.0$ $\mathbf{54.0}$$\pm 9.0$ $23.5$$\pm 19.5$ $38.5$$\pm 1.5$ $39.5$$\pm 11.5$ $20.5$$\pm 20.5$ $339.0$
(4) What makes model-based offline RL work in AntMaze?
LEQ shows outstanding performance compared to previous offline model-based RL methods, especially in large and ultra mazes. To understand which aspects of LEQ enabled this success, we applied its changes to MOBILE and analyzed the impact. The results are detailed in Table 6.
Table 6: Impact of hyperparameters in MOBILE on AntMaze. MOBILE uses the hyperparameters from MOBILE: $\beta=0.95$, $\gamma=0.99$, and $R=5$, whereas LEQ uses $\beta=0.25$, $\gamma=0.997$, and $R=10$. The results show that $\beta$ is the most critical hyperparameter that makes MOBILE work in AntMaze.
  Hyperparams. umaze medium large ultra Total
$\beta$ $\gamma$ $R$ umaze diverse play diverse play diverse play diverse
We first re-implement MOBILE with some technical tricks used in LEQ (denoted as MOBILE): LayerNorm (Ba et al., 2016), SymLog (Hafner et al., 2023), single Q-network, and no target Q-value clipping. However, MOBILE still achieves a barely non-zero score, $14.0$.
@@
We found that reducing the ratio $\beta$ between the losses from imaginary rollouts and dataset transitions is key to make MOBILE work (i.e. achieving meaningful performances in umaze and medium mazes, with a total score of $213.0$). This adjustment also allows for a higher discount rate and longer imagination horizon, yielding the best results for MOBILE. We suggest that utilizing the true transition from the dataset is important in long-horizon tasks, which has been undervalued in prior works. We provide additional extensive ablation results on LEQ in Appendix D.
(5) Imagination length $H$ and dataset expansion.
@@
As shown in Table 7, the performance increases when it goes to $H=10$ from $H=5$, but it drops when $H=15$. This result shows the trade-off of using the world model – the further the agent imagines, the more the agent becomes robust to the error of the critic, but the more it becomes prone to the error from the model prediction.
@@
We also evaluate LEQ without the dataset expansion. In AntMaze, the results with and without the dataset expansion are similar, as shown in Table 7. On the other hand, the dataset expansion makes the policy more stable and better in the D4RL MuJoCo tasks (in Appendix, Table 20).
@@
Table 7: LEQ with different imagination length $H$ and data expansion. A longer $H$ can mitigate critic biases, while increasing model errors, which leads to poor performance. Each number is averaged over $5$ random seeds.
Dataset $\mathbf{H=10}$ (ours) $H=5$ $H=15$ w/o dataset expansion
@@
6 Limitations
Following prior work on model-based offline RL (Sun et al., 2023; Jeong et al., 2023), we assume access to the ground-truth termination function of a task, different from online model-based RL approaches, which learn a termination function from interactions. As shown in Table 22, Using a learned termination function instead of the ground-truth termination function results in a significant performance drop ($461.8\rightarrow 232.6$), particularly in diverse datasets ($233.2\rightarrow 66.8$), where termination signals are limited because the dataset is collected by navigating to randomly selected goals. While relying on the ground-truth termination function simplifies the problem, it limits the applicability of the method to scenarios where this information is readily available. Extending the proposed approach to learn terminal signals from the dataset would be an immediate next step.
@@ 7 Conclusion
In this paper, we propose a novel offline model-based reinforcement learning method, LEQ, which uses expectile regression to get a conservative evaluation of a policy from model-generated trajectories. Expectile regression eases the pain of constructing the whole distribution of Q-targets and allows for learning a conservative Q-function via sampling. Combined with $\lambda$-returns in both critic and policy updates for the imaginary rollouts, the policy can receive learning signals that are more robust to both model errors and critic errors. We empirically show that LEQ robustly improves the performance of model-based approaches in various domains, including state-based locomotion, long-horizon navigation, and visual control tasks.
Acknowledgments
We would like to thank Junik Bae for helpful discussion. This work was supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant (RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)) and the National Research Foundation of Korea (NRF) grant (RS-2024-00333634) funded by the Korean Government (MSIT). Kwanyoung Park was supported by the Electronics and Telecommunications Research Institute (ETRI) grant (24ZR1100) and IITP grant (RS-2024-00509279, Global AI Frontier Lab) funded by the Korea government (MSIT).
Ethics Statement
Our method aims to increase the ability of autonomous agents, such as robots and self-driving cars, to learn from static, offline data without interacting with the world. This enables autonomous agents to utilize data with diverse qualities (not necessarily from experts). We believe that this paper does not have any immediate negative ethical concerns.
Reproducibility Statement
To ensure the reproducibility of our work, we provide the full code of LEQ in the supplementary materials, along with instructions to replicate the experiments presented in the paper. We provide the experimental details in Appendix A and the proof on the derivation of our surrogate policy objective in Appendix B.
@@ Appendix A Training Details
All experiments are done on a single RTX 4090 GPU and $8$ AMD EPYC 9354 CPU cores. For state-based environments, we use $5$ different random seeds for each experiment and report the mean and standard deviation, while each offline RL experiment takes $2$ hours for ours, $12$ hours for MOBILE, and $24$ hours for CBOP. For pixel-based environments, we use $3$ different random seeds and each experiment takes $2$ hours for ours and $8$ hours for ROSMO.
@@
For state-based locomotion tasks, we use the datasets provided by D4RL [Fu et al., 2020] and NeoRL [Qin et al., 2022]. Following IQL [Kostrikov et al., 2022], we normalize rewards using the maximum and minimum returns of all trajectories. We use the true termination functions of the environments, implemented in MOBILE [Sun et al., 2023]. For pixel-based environments, we do not normalize the rewards.
For AntMaze tasks, we use the datasets provided by D4RL [Fu et al., 2020]. Following IQL [Kostrikov et al., 2022], we subtract $1$ from the rewards in the datasets so that the agent receives $-1$ for each step and 00 on termination. We use the true termination functions of the environments. The termination functions of the AntMaze tasks are not deterministic because a goal of a maze is randomized every time the environment is reset. Nevertheless, we follow the implementation of CBOP [Jeong et al., 2023], where the termination region is set to a circle around the mean of the goal distribution with the radius $0.5$.
Implementation details of compared methods.
For all compared methods, we use the results from their corresponding papers when available. For IQL [Kostrikov et al., 2022], we run the official implementation with $5$ seeds to reproduce the results for the random datasets in D4RL and NeoRL. For the AntMaze tasks, we run the official implementation of MOBILE and CBOP with $5$ random seeds. Please note that the original MOBILE implementation does not use the true termination function, so we replace it with our termination function. For MOPO, COMBO, and RAMBO, we use the results reported in RAMBO [Rigter et al., 2022]. For DMControl tasks, we replace the categorical distribution of the policy with gaussian distribution of the official ROSMO codebase, and run the experiments with sampling hyperparameter $N=4$.
@@
For state-based environments, we use the architecture and training script from OfflineRL-Kit [Sun, 2023], matching the implementation of MOBILE [Sun et al., 2023]. Each world model is implemented as a $4$-layer MLPs with the hidden layer size of $200$. We construct an ensemble of world models by selecting $5$ out of $7$ models with the best validation scores. We pretrain the ensemble of world models for each of $5$ random seeds (i.e. training in total $35$ world models and using $25$ models), which takes approximately $5$ hours in average. For pixel-based environments, we use the 12M model of DreamerV3 [Hafner et al., 2023] and pretrain the world model using its loss function. We follow the implementation of OfflineDV2 [Lu et al., 2023], training 7 ensemble of world models for stochastic latent prediction, which takes $4$ hours in average. We select one model for each step, following the design choice of MOBILE, as we found that randomly choosing a model every step can make imaginary rollouts more robust to model biases, leading to better performance.
@@
For state-based environments, we use $3$-layer MLPs with size of $256$ both for the policy network and the critic network. We use layer normalization [Ba et al., 2016] to prevent catastrophic over/underestimation [Ball et al., 2023], and squash the state inputs using symlog to keep training stable from outliers in long-horizon model rollouts [Hafner et al., 2023]. For pixel-based environments, we use the architecture of the 12M model in DreamerV3 models.
Pretraining policy and critic networks.
For some environments, we found that a randomly initialized policy can lead to abnormal rewards or transition prediction from the world models in the early stage, leading to unstable training [Jelley et al., 2024]. Following CBOP [Jeong et al., 2023], we pretrain a policy $\pi_{\theta}$ and a critic $Q_{\phi}$ using behavioral cloning and FQE [Le et al., 2019], respectively for state-based experiments. We use a slightly different implementation of FQE from the original implementation, where the $\arg\min$ operation is approximated with mini-batch gradient descent, similar to standard Q-learning as shown in Algorithm 2.
Algorithm 2 FQE: Fitted Q Evaluation [Le et al., 2019]
@@
4:     $y_{i}=\texttt{sg}(r_{i}+Q_{\phi}(\mathbf{s}^{\prime}_{i},\pi_{\theta}(\mathbf{s}^{\prime}_{i})))$triangleright $\texttt{sg}(\cdot)$ is stop-gradient operator
@@
Table 8: Comparisons with the prior model-based methods and the baseline method. The hyperparameters same with LEQ are colored in blue; others are colored in red.
Components CBOP MOBILE MOBILE LEQ (ours)
Training scheme MVE [Feinberg et al., 2018] MBPO [Janner et al., 2019] MBPO [Janner et al., 2019] Dreamer [Hafner et al., 2023]
Conservatism Lower-confidence bound Lower-confidence bound Lower-confidence bound Lower expectile
Policy Stochastic Stochastic Stochastic Deterministic
Policy objective $Q(\mathbf{s},\mathbf{a})$ $Q(\mathbf{s},\mathbf{a})$ $Q(\mathbf{s},\mathbf{a})$ $\lambda$-returns
Policy pretraining BC BC
# of critics 20-50 2 1 1
Critic objective Multi-step (adaptive weighting) One-step One-step $\lambda$-returns
Critic pretraining FQE [Le et al., 2019] FQE [Le et al., 2019]
Horizon length ($H$) 10 1 1 10
Expansion length ($R$) 1 or 5 10 5
Discount rate ($\gamma$) 0.99 0.99 0.997 0.997
$\beta$ in Equation 7 1.0 0.95 0.25 0.25
Impl. tricks Clip Q-values with 00 LayerNorm + Symlog LayerNorm + Symlog
Running time 24h 12h 40m 2h